1. In early 2008 an investor purchased and remodeled a hotel to handle projected increases in local area demand. By the end of the year, it was clear that the recession would cut this demand significantly. Under these new demand conditions the hotel will be able to sell 20,000 room nights that will cost on average $50 per room night to service. The investor spent $20 million on remodeling and has a cost of capital of 10%. What is the break-even point for the hotel?
Answer #1

The Total Cost of investment after 1 year =20,000,000*(1+cost of capital)=20000000*1.1 =22,000,000
Sales Cost -Variable Cost =Total cost/Number of rooms =22,000,000/20000
Breakeven Sales Cost =22,000,000/20000+50 =1150